# Aeon: High-Performance Neuro-Symbolic Memory Management for Long-Horizon LLM Agents

### v1

@@ Abstract
Large Language Models (LLMs) are fundamentally constrained by the quadratic computational cost of self-attention and the “Lost in the Middle” phenomenon, where reasoning capabilities degrade as context windows expand. Existing solutions, primarily “Flat RAG” architectures relying on vector databases, treat memory as an unstructured bag of embeddings. This approach fails to capture the hierarchical and temporal structure of long-horizon interactions, leading to “Vector Haze”—the retrieval of disjointed facts lacking episodic continuity. We propose Aeon, a Neuro-Symbolic Cognitive Operating System that redefines memory not as a static store, but as a managed OS resource. Aeon structures memory into a Memory Palace (a spatial index implemented via Atlas, a SIMD-accelerated Page-Clustered Vector Index that combines small-world graph navigation with B+ Tree-style disk locality to minimize read amplification) and a Trace (a neuro-symbolic episodic graph). We introduce the Semantic Lookaside Buffer (SLB), a predictive caching mechanism that exploits conversational locality to achieve sub-millisecond retrieval latencies. Benchmarks demonstrate that Aeon achieves ms retrieval latency on conversational workloads while ensuring state consistency via a zero-copy C++/Python bridge, effectively enabling persistent, structured memory for autonomous agents.
@@ 1 Introduction
The rapid evolution of Large Language Models (LLMs) has been defined by a relentless scaling of parameters and training data, yet the fundamental architecture remains bound by the Context Bottleneck. The Transformer’s self-attention mechanism, while transformative, imposes a quadratic time and space complexity, , relative to the input sequence length. Although recent optimization techniques—such as sparse attention, RingAttention, and hardware-aware kernel fusion—have theoretically extended context windows to 1 million tokens and beyond, the utility of this context does not scale linearly. Empirical evidence highlights a distinct degradation in reasoning capabilities over these extended horizons, a phenomenon widely characterized as being “Lost in the Middle” [8]. When relevant information is buried in the center of a massive context window, model performance deteriorates often below the baseline of closed-book inference. As autonomous agents are tasked with increasingly complex, long-horizon objectives spanning days or weeks, the reliance on these transient, volatile context windows becomes untenable. The model cannot simply attend to all of history; it must select what is potentially relevant before attention is even applied. However, the selection mechanism itself is often limited by the same architectural constraints it seeks to mitigate. We argue that the simplistic view of context as a sliding window is insufficient for general-purpose autonomy. Instead, we require a memory architecture that is as persistent, structured, and deterministically managed as the storage hierarchies found in classical operating systems.
@@
The prevailing industry response to the context limitation has been the widespread adoption of Retrieval-Augmented Generation (RAG). In its most common form, “Flat RAG,” this approach offloads information preservation to vector databases that perform Approximate Nearest Neighbor (ANN) search over unstructured lists of embeddings. Technologies such as HNSW (Hierarchical Navigable Small World) graphs or inverted file indices allow for efficient retrieval of semantically similar chunks. However, while effective for simple, one-shot question-answering tasks, Flat RAG fails to model the structure of extended interaction. It treats memory as a featureless plane—a “bag of vectors”—where the temporal evolution of a conversation, the causal lineage of decisions, and the hierarchical relationship between concepts are lost. We term this failure mode “Vector Haze”: the retrieval of semantically similar but episodically disjointed facts that confuse rather than aid the agent. A Flat RAG system has no notion of “where” it is in a conversation; it only knows “what” matches the current query vector in high-dimensional space. It lacks the ability to backtrack to a previous state, to branch a conversation into potential futures, or to understand the narrative arc that led to the current moment. For an agent to maintain coherence over thousands of turns, it requires more than a semantic search engine; it requires a state machine that understands the topology of its own experience.
@@
We propose a paradigm shift from treating memory as a passive database retrieval problem to treating it as an active resource management problem within a Cognitive Operating System. In this view, memory management becomes analogous to virtual memory management in traditional OS kernels. We introduce Aeon, a system that formalizes these operations. Allocation corresponds to the deliberate writing of new semantic concepts into a structured Atlas; paging transforms into the loading of relevant semantic clusters into a Semantic Lookaside Buffer (SLB) for immediate, low-latency access; and context switching is re-framed as the deterministic movement between branches of a decision tree. By enforcing these rigorous abstractions, Aeon transforms the probabilistic chaos of vector search into a deterministic navigational process. This empowers the agent to utilize a “Memory Palace”—a spatial index where information is stored not just by what it means, but by where it belongs relative to other concepts in the agent’s ontology. This spatial locality enables the system to predictively pre-fetch memory, much like a CPU pre-fetches instructions based on spatial and temporal locality, drastically reducing the latency penalty of retrieval and allowing the agent to "think" at the speed of conversation.
@@
Our contributions in this paper are threefold. First, we introduce Atlas, a high-performance, memory-mapped B+ Tree that organizes uniform vectors into a navigable, hierarchical index. Unlike HNSW or IVFPQ indices which optimize purely for recall at the expense of insert performance and structure, Atlas optimizes for semantic locality and stable modification. It utilizes a custom SIMD-accelerated math kernel to perform metric comparisons, ensuring that the tree structure strictly adheres to the geometry of the embedding space. Second, we present the Trace, a neuro-symbolic directed acyclic graph (DAG) that explicitly tracks the agent’s episodic state. The Trace records the traversal path of the agent, creating distinct nodes for User inputs, System responses, and intermediate thoughts, linked by typed edges (CAUSAL, NEXT, REFERS_TO). This enables capabilities such as backtracking and context anchoring that are impossible in flat vector stores. Finally, we demonstrate the implementation of a zero-copy architecture bridging C++23 and Python. By harnessing the nanobind library, Aeon exposes internal C++ memory structures as read-only NumPy arrays, eliminating serialization overhead. This allows Aeon to achieve retrieval latencies of ms on standard conversational workloads, validating the feasibility of this Cognitive OS approach for real-time, interactive agents that require both the speed of systems programming and the flexibility of high-level reasoning.
@@ 2 Related Work
We position Aeon within the broader landscape of neural memory systems, contrasting its Cognitive Operating System architecture against existing approaches in retrieval, memory management, and neuro-symbolic reasoning. While individual components of Aeon have precedents in isolation, we argue that the lack of a unified, high-performance kernel has limited the emergence of truly long-horizon agents.
@@ 2.1 Retrieval-Augmented Generation (RAG)
The dominant paradigm for grounding Large Language Models (LLMs) is Retrieval-Augmented Generation (RAG), typically implemented using Dense Passage Retrieval (DPR) [7] and Approximate Nearest Neighbor (ANN) search indices like FAISS [5] or HNSW [9]. These systems rely on what we characterize as "Flat RAG": a single, monolithic vector space where every query is treated as an independent event, disconnected from temporal or causal context.
@@ 2.2 Memory-Augmented LLMs
However, MemGPT is a logical framework rather than a systems-level implementation. It functions in "User Space" (typically Python), relying on the LLM itself to manage memory calls via prompt engineering. This introduces significant latency and leaves the physical layout of memory unmanaged. Aeon moves this responsibility to "The Kernel," implementing memory management in high-performance C++23. By treating memory operations as low-level system calls, Aeon achieves sub-millisecond retrieval latencies that are impossible with prompt-based management.
@@ 2.4 Operating System Primitives for AI
Finally, the concept of an "AI Kernel" has been proposed in various forms, often serving as a metaphor for cloud orchestration layers or "Glue Code" libraries like LangChain [1]. While these frameworks provide essential developer tooling, they do not function as operating systems in the traditional sense; they do not manage physical memory layout, thread scheduling, or cache coherency.
Aeon distinguishes itself by implementing true OS primitives tailored for semantic data. The Semantic Lookaside Buffer (SLB) is a direct parallel to the Translation Lookaside Buffer (TLB) in CPU architecture. Instead of just caching raw vectors, the SLB predicts and prefetches semantic clusters relevant to the current thread of thought. This moves the paradigm from reactive retrieval to proactive memory management, solving the "Context Bottleneck" at the architectural level.
@@ 3.1 Design Philosophy: The Core-Shell Model
The Core (Ring 0): Implemented in C++23, the Core is responsible for all high-frequency, low-latency operations. This includes vector similarity search, tree traversal, and memory management. It operates directly on raw memory pages and leverages hardware acceleration (SIMD AVX-512) to maximize throughput.
The Shell (Ring 3): Implemented in Python 3.12, the Shell manages high-level control logic, including Large Language Model (LLM) interaction, prompt engineering, and graph topology management. It serves as the "cognitive" layer that orchestrates the system’s behavior.
@@ 3.2.2 Storage and Access
The Atlas resides entirely on persistent storage (NVMe SSD) but is mapped into the process’s virtual address space using the POSIX mmap system call. This allows the OS virtual memory subsystem to handle page caching transparently. We explicitly avoid standard C++ heap allocations (e.g., new, malloc) for node data to ensure data contiguity and cache locality.
@@ 3.2.3 Greedy SIMD Descent
The system selects the next node and recurses until a local optimum (leaf) is reached. The complexity of this descent is , where is the effective branching factor and is the total number of nodes in the Atlas. All vector operations are vectorized using AVX-512 intrinsics, allowing 16 floating-point operations per cycle.
@@ 3.3.3 Navigation
The Trace enables the LLM to perform "backtracking." By traversing (inverse temporal edges), the agent can effectively "rewind" its cognitive state to a previous turn to resolve ambiguities or correct context drift.
@@ 3.4 The Zero-Copy Interface
To strictly enforce the Zero-Copy Constraint, we utilize nanobind to expose C++ memory structures to Python. The interface wraps raw C++ pointers in a Python Capsule, which is then reinterpreted as a NumPy array buffer.
@@ 4 The Semantic Lookaside Buffer
The core contribution of Aeon is the Semantic Lookaside Buffer (SLB), a high-performance caching mechanism designed to alleviate the latency overhead of traversing massive high-dimensional indices. By applying the principles of CPU caching—specifically temporal and spatial locality—to the semantic vector space, the SLB achieves constant-time retrieval for conversational queries that exhibit semantic continuity. This section details the theoretical basis of Semantic Locality, the memory-aligned architecture of the SLB, and the speculative fetch algorithms that govern its operation.
@@ 4.1 Theory: Semantic Locality
Traditional caching strategies, such as Least Recently Used (LRU) or Least Frequently Used (LFU), rely on address transparency and the assumption that repeated access to the exact same memory address is common. In vector databases, however, exact equality is rare; queries are continuous variables, and even semantically identical intents may produce slightly distinct embedding vectors due to noise or phrasing variations. Thus, we introduce the concept of Semantic Inertia.
@@
Formally, let be the high-dimensional semantic space and be the query vector at step . We posit that the conditional probability of the distance between subsequent queries being within a small locality radius is significantly non-zero:
@@
This implies that for a significant subset of queries, the optimal search starting point is not the global Root of the hierarchical navigable small world (HNSW) graph or the Atlas tree, but rather the result of the previous query. If the semantic delta is sufficiently small, the target node for is likely the same node as , or one of its immediate semantic neighbors. By caching these recent access points, we can bypass the tree traversal entirely.
@@ 4.2 Architecture of the SLB
Search Strategy: Brute-Force SIMD. Crucially, we do not use an approximate nearest neighbor index for the SLB itself. Because is small (), we can perform an exhaustive linear scan of all entries using AVX-512 instructions. The cost of this operation is , but due to perfect hardware prefetching and zero pointer chasing, the wall-clock time is significantly lower than even a few steps of an tree traversal. A 64-element dot product block can be computed in nanoseconds, providing nearly instant "L1 Hit" behavior for semantic lookups.
@@ 4.3 The Speculative Fetch Algorithm
Update: The result of the query (whether from a hit or a miss-then-retrieval) is inserted into the SLB. We employ a simplified Least Recently Used (LRU) eviction policy, implemented via the ring buffer pointer or a lightweight timestamp, to ensure fresh context is prioritized.
@@ 4.4 Predictive Prefetching
While the current implementation focuses on reducing lookup latency, the SLB architecture enables powerful future optimization such as Async Prefetching. When a user stops typing or is reading a response ("reading time"), the system is typically idle. We can exploit this downtime. Upon a successful SLB hit, the system can speculatively load the children of the hit node from the main memory-mapped file into the SLB.
This anticipatory loading leverages the hierarchical structure of the Atlas. If a user is querying a specific sub-topic (e.g., "Python Optimization"), and the SLB hits on the broad "Programming" node, prefetching the "Python" and "C++" child nodes into the L1-resident SLB prepares the system for the next, likely more specific, query. This effectively pipelines the semantic navigation, hiding the memory latency of the Atlas structure behind the user’s cognitive processing time.
@@ 5 Experimental Methodology
This section describes the experimental setup used to evaluate Aeon’s performance characteristics. All experiments were conducted five times, and we report the median value along with the 25th and 75th percentiles to account for variance.

### v2

@@ Abstract
Large Language Models (LLMs) are fundamentally constrained by the quadratic computational cost of self-attention and the “Lost in the Middle” phenomenon, where reasoning capabilities degrade as context windows expand. Existing solutions, primarily “Flat RAG” architectures relying on vector databases, treat memory as an unstructured bag of embeddings. This approach fails to capture the hierarchical and temporal structure of long-horizon interactions, leading to “Vector Haze”: the retrieval of disjointed facts lacking episodic continuity. This paper proposes Aeon, a Neuro-Symbolic Cognitive Operating System that redefines memory not as a static store, but as a managed OS resource. Aeon structures memory into a Memory Palace (a spatial index implemented via Atlas, a SIMD-accelerated Page-Clustered Vector Index that combines small-world graph navigation with B+ Tree-style disk locality to minimize read amplification) and a Trace (a neuro-symbolic episodic graph). The Semantic Lookaside Buffer (SLB), a predictive caching mechanism, exploits conversational locality to achieve sub-millisecond retrieval latencies. Benchmarks on Apple M4 Max demonstrate that Aeon achieves s effective retrieval latency on conversational workloads (with 85%+ SLB hit rates), while ensuring state consistency via a sub-microsecond zero-copy C++/Python bridge (334ns for 10MB payloads), effectively enabling persistent, structured memory for autonomous agents.
@@ 1 Introduction
The rapid evolution of Large Language Models (LLMs) has been defined by a relentless scaling of parameters and training data, yet the fundamental architecture remains bound by the Context Bottleneck. The Transformer’s self-attention mechanism, while transformative, imposes a quadratic time and space complexity, , relative to the input sequence length. Although recent optimization techniques (sparse attention, RingAttention, and hardware-aware kernel fusion) have theoretically extended context windows to 1 million tokens and beyond, the utility of this context does not scale linearly. Empirical evidence highlights a distinct degradation in reasoning capabilities over these extended horizons, a phenomenon widely characterized as being “Lost in the Middle” [8]. When relevant information is buried in the center of a massive context window, model performance deteriorates, often below the baseline of closed-book inference. As autonomous agents are tasked with increasingly complex, long-horizon objectives spanning days or weeks, the reliance on these transient, volatile context windows becomes untenable. The model cannot simply attend to all of history; it must select what is potentially relevant before attention is even applied. However, the selection mechanism itself is often limited by the same architectural constraints it seeks to mitigate. The simplistic view of context as a sliding window is insufficient for general-purpose autonomy. Instead, what is required is a memory architecture that is as persistent, structured, and deterministically managed as the storage hierarchies found in classical operating systems.
@@
The prevailing industry response to the context limitation has been the widespread adoption of Retrieval-Augmented Generation (RAG). In its most common form, “Flat RAG,” this approach offloads information preservation to vector databases that perform Approximate Nearest Neighbor (ANN) search over unstructured lists of embeddings. Technologies such as HNSW (Hierarchical Navigable Small World) graphs or inverted file indices allow for efficient retrieval of semantically similar chunks. However, while effective for simple, one-shot question-answering tasks, Flat RAG fails to model the structure of extended interaction. It treats memory as a featureless plane (a “bag of vectors”) where the temporal evolution of a conversation, the causal lineage of decisions, and the hierarchical relationship between concepts are lost. This failure mode is termed “Vector Haze”: the retrieval of semantically similar but episodically disjointed facts that confuse rather than aid the agent. A Flat RAG system has no notion of “where” it is in a conversation; it only knows “what” matches the current query vector in high-dimensional space. It lacks the ability to backtrack to a previous state, to branch a conversation into potential futures, or to understand the narrative arc that led to the current moment. For an agent to maintain coherence over thousands of turns, it requires more than a semantic search engine; it requires a state machine that understands the topology of its own experience.
@@
This paper proposes a paradigm shift from treating memory as a passive database retrieval problem to treating it as an active resource management problem within a Cognitive Operating System. In this view, memory management becomes analogous to virtual memory management in traditional OS kernels. Aeon formalizes these operations. Allocation corresponds to the deliberate writing of new semantic concepts into a structured Atlas; paging transforms into the loading of relevant semantic clusters into a Semantic Lookaside Buffer (SLB) for immediate, low-latency access; and context switching is re-framed as the deterministic movement between branches of a decision tree. By enforcing these rigorous abstractions, Aeon transforms the probabilistic chaos of vector search into a deterministic navigational process. This empowers the agent to utilize a “Memory Palace,” a spatial index where information is stored not just by what it means, but by where it belongs relative to other concepts in the agent’s ontology. This spatial locality enables the system to predictively pre-fetch memory, much like a CPU pre-fetches instructions based on spatial and temporal locality, drastically reducing the latency penalty of retrieval and allowing the agent to “think” at the speed of conversation.
@@
The contributions of this paper are threefold. First, it introduces Atlas, a high-performance, memory-mapped B+ Tree that organizes uniform vectors into a navigable, hierarchical index. Unlike HNSW or IVFPQ indices which optimize purely for recall at the expense of insert performance and structure, Atlas optimizes for semantic locality and stable modification. It utilizes a custom SIMD-accelerated math kernel to perform metric comparisons, ensuring that the tree structure strictly adheres to the geometry of the embedding space. Second, the paper presents the Trace, a neuro-symbolic directed acyclic graph (DAG) that explicitly tracks the agent’s episodic state. The Trace records the traversal path of the agent, creating distinct nodes for User inputs, System responses, and intermediate thoughts, linked by typed edges (CAUSAL, NEXT, REFERS_TO). This enables capabilities such as backtracking and context anchoring that are impossible in flat vector stores. Finally, the paper demonstrates the implementation of a zero-copy architecture bridging C++23 and Python. By harnessing the nanobind library, Aeon exposes internal C++ memory structures as read-only NumPy arrays, eliminating serialization overhead. This allows Aeon to achieve effective retrieval latencies of s on standard conversational workloads, validating the feasibility of this Cognitive OS approach for real-time, interactive agents that require both the speed of systems programming and the flexibility of high-level reasoning.
@@ 2 Related Work
Aeon is positioned within the broader landscape of neural memory systems, contrasting its Cognitive Operating System architecture against existing approaches in retrieval, memory management, and neuro-symbolic reasoning. While individual components of Aeon have precedents in isolation, the lack of a unified, high-performance kernel has limited the emergence of truly long-horizon agents.
@@ 2.1 Retrieval-Augmented Generation (RAG)
The dominant paradigm for grounding Large Language Models (LLMs) is Retrieval-Augmented Generation (RAG), typically implemented using Dense Passage Retrieval (DPR) [7] and Approximate Nearest Neighbor (ANN) search indices like FAISS [5] or HNSW [9]. These systems rely on what is characterized here as “Flat RAG”: a single, monolithic vector space where every query is treated as an independent event, disconnected from temporal or causal context.
@@ 2.2 Memory-Augmented LLMs
However, MemGPT is a logical framework rather than a systems-level implementation. It functions in “User Space” (typically Python), relying on the LLM itself to manage memory calls via prompt engineering. This introduces significant latency and leaves the physical layout of memory unmanaged. Aeon moves this responsibility to “The Kernel,” implementing memory management in high-performance C++23. By treating memory operations as low-level system calls, Aeon achieves sub-microsecond retrieval latencies that are impossible with prompt-based management.
@@ 2.4 Operating System Primitives for AI
Finally, the concept of an “AI Kernel” has been proposed in various forms, often serving as a metaphor for cloud orchestration layers or “Glue Code” libraries like LangChain [1]. While these frameworks provide essential developer tooling, they do not function as operating systems in the traditional sense; they do not manage physical memory layout, thread scheduling, or cache coherency.
Aeon distinguishes itself by implementing true OS primitives tailored for semantic data. The Semantic Lookaside Buffer (SLB) is a direct parallel to the Translation Lookaside Buffer (TLB) in CPU architecture. Instead of just caching raw vectors, the SLB predicts and prefetches semantic clusters relevant to the current thread of thought. This moves the paradigm from reactive retrieval to proactive memory management, solving the “Context Bottleneck” at the architectural level.
@@ 3.1 Design Philosophy: The Core-Shell Model
The Core (Ring 0): Implemented in C++23, the Core is responsible for all high-frequency, low-latency operations. This includes vector similarity search, tree traversal, and memory management. It operates directly on raw memory pages and leverages hardware acceleration (SIMD AVX-512 via SIMDe) to maximize throughput.
The Shell (Ring 3): Implemented in Python 3.12, the Shell manages high-level control logic, including Large Language Model (LLM) interaction, prompt engineering, and graph topology management. It serves as the “cognitive” layer that orchestrates the system’s behavior.
@@ 3.2.2 Storage and Access
The Atlas resides entirely on persistent storage (NVMe SSD) but is mapped into the process’s virtual address space using the POSIX mmap system call. This allows the OS virtual memory subsystem to handle page caching transparently. Standard C++ heap allocations (e.g., new, malloc) are explicitly avoided for node data to ensure data contiguity and cache locality.
@@ 3.2.3 Greedy SIMD Descent
The system selects the next node and recurses until a local optimum (leaf) is reached. The complexity of this descent is , where is the effective branching factor and is the total number of nodes in the Atlas. All vector operations are vectorized using AVX-512 intrinsics via the SIMDe portability layer, enabling 16 floating-point operations per cycle on x86-64 and equivalent throughput on ARM64 via NEON translation.
@@ 3.3.3 Navigation
The Trace enables the LLM to perform “backtracking.” By traversing (inverse temporal edges), the agent can effectively “rewind” its cognitive state to a previous turn to resolve ambiguities or correct context drift.
@@ 3.4 The Zero-Copy Interface
To strictly enforce the Zero-Copy Constraint, Aeon utilizes nanobind to expose C++ memory structures to Python. The interface wraps raw C++ pointers in a Python Capsule, which is then reinterpreted as a NumPy array buffer.
@@ 4 The Semantic Lookaside Buffer
The core contribution of Aeon is the Semantic Lookaside Buffer (SLB), a high-performance caching mechanism designed to alleviate the latency overhead of traversing massive high-dimensional indices. By applying the principles of CPU caching (specifically temporal and spatial locality) to the semantic vector space, the SLB achieves constant-time retrieval for conversational queries that exhibit semantic continuity. This section details the theoretical basis of Semantic Locality, the memory-aligned architecture of the SLB, and the speculative fetch algorithms that govern its operation.
@@ 4.1 Theory: Semantic Locality
Traditional caching strategies, such as Least Recently Used (LRU) or Least Frequently Used (LFU), rely on address transparency and the assumption that repeated access to the exact same memory address is common. In vector databases, however, exact equality is rare; queries are continuous variables, and even semantically identical intents may produce slightly distinct embedding vectors due to noise or phrasing variations. Thus, the concept of Semantic Inertia is introduced.
@@
Formally, let be the high-dimensional semantic space and be the query vector at step . The following is posited: the conditional probability of the distance between subsequent queries being within a small locality radius is significantly non-zero:
@@
This implies that for a significant subset of queries, the optimal search starting point is not the global Root of the hierarchical navigable small world (HNSW) graph or the Atlas tree, but rather the result of the previous query. If the semantic delta is sufficiently small, the target node for is likely the same node as , or one of its immediate semantic neighbors. By caching these recent access points, the tree traversal can be bypassed entirely.
@@ 4.2 Architecture of the SLB
Search Strategy: Brute-Force SIMD. Crucially, an approximate nearest neighbor index is not used for the SLB itself. Because is small (), an exhaustive linear scan of all entries can be performed using AVX-512 instructions. The cost of this operation is , but due to perfect hardware prefetching and zero pointer chasing, the wall-clock time is significantly lower than even a few steps of an tree traversal. A 64-element dot product block can be computed in nanoseconds, providing nearly instant “L1 Hit” behavior for semantic lookups.
@@ 4.3 The Speculative Fetch Algorithm
Update: The result of the query (whether from a hit or a miss-then-retrieval) is inserted into the SLB. A simplified Least Recently Used (LRU) eviction policy, implemented via the ring buffer pointer or a lightweight timestamp, ensures fresh context is prioritized.
@@ 4.4 Predictive Prefetching
While the current implementation focuses on reducing lookup latency, the SLB architecture enables powerful future optimization such as Async Prefetching. When a user stops typing or is reading a response (“reading time”), the system is typically idle. This downtime can be exploited. Upon a successful SLB hit, the system can speculatively load the children of the hit node from the main memory-mapped file into the SLB.
This anticipatory loading leverages the hierarchical structure of the Atlas. If a user is querying a specific sub-topic (e.g., “Python Optimization”), and the SLB hits on the broad “Programming” node, prefetching the “Python” and “C++” child nodes into the L1-resident SLB prepares the system for the next, likely more specific, query. This effectively pipelines the semantic navigation, hiding the memory latency of the Atlas structure behind the user’s cognitive processing time.
@@ 5 Experimental Methodology
This section describes the experimental setup used to evaluate Aeon’s performance characteristics. All experiments were conducted five times, and the median value is reported along with the 25th and 75th percentiles to account for variance.